The personnel department of a large corporation wants to estimate the family dental expenses of its employees to determine the feasibility of providing a dental insurance plan. A random sample of six employees reveals the following family dental expenses: $180, $260, $60, $40, $100, and $80. It is known that dental expenses follow a normal distribution. Construct a 90% confidence interval for the average family dental expenses for all employees of this corporation.
Answer #1

From the given sample data,

\bar{x} = \sum Xi / n

= 120

S = sqrt( \sum Xi2 - n\bar{x}2 / n-1)

= 83.9047

90% confidence interval for \mu is

\bar{x} - t\alpha/2 * S / sqrt(n) < \mu < \bar{x} + t\alpha/2 * S / sqrt(n)

120 - 2.015 * 83.9047 / sqrt(6) < \mu < 120 + 2.015 * 83.9047 / sqrt(6)

50.9783 < \mu < 189.0217

90% CI for \mu is (50.9783 , 189.0217)


answered by: ANURANJAN SARSAM